# Cosmology with voids

### v3

@@ 6.1 Reconstruction methods and peculiar velocity surveys
The large-scale structure revealed in redshift surveys, though affected by line-of-sight peculiar motion of individual galaxies, provides information about the underlying mass distribution in real space. If this real-space distribution were known, then linear theory would predict the peculiar velocities and hence the redshift-space location of galaxies within it. This mapping, at least on scales above $O(10)$ $h^{-1}$pc where the theory applies, is invertible. The real-space mass distribution and peculiar velocity flows can be reconstructed from redshift data (Yahil et al., 1991; Dekel & Rees, 1994; Fisher et al., 1994; Nusser & Davis, 1994; Gramann et al., 1994; Tegmark & Bromley, 1995; Zaroubi & Hoffman, 1996; Taylor & Hamilton, 1996; Hamilton, 1998; Taylor & Valentine, 1999; Burkey & Taylor, 2004; Iršič & Slosar, 2011; Koda et al., 2014; Davis & Scrimgeour, 2014; White, 2014; Watkins & Feldman, 2015; Carrick et al., 2015; Howlett et al., 2017a, b; Adams & Blake, 2020; Said et al., 2020; Whitford et al., 2022). Recent work has extended reconstruction methods to include smaller scales where nonlinear dynamics affect redshift-space distortions (Hamaus et al., 2017; Ganeshaiah Veena et al., 2023; Qin et al., 2023; Chen et al., 2023, 2024; Lilow et al., 2024; Wang & Yang, 2024).
Density reconstruction methods have been designed for the study of voids (Paz et al., 2013; Nadathur et al., 2019a). Because the details of void selection are critical in measures of the void-galaxy correlator, reconstruction of the real-space density field prior to void identification is important so that redshift-space distortions do not introduce adverse selection effects (Chuang et al., 2017; Nadathur et al., 2020a).
Reconstruction methods typically depend on assumptions about the background cosmology (Kaiser, 1987; Dekel & Rees, 1994) — the Hubble expansion rate $H$ and velocity growth factor $f$ are often required parameters in a reconstruction application. Therefore, reconstructed density and velocity fields on their own are not well suited for parameter estimation. They return the cosmological information from which they were derived. Following the lead of (Dekel & Rees, 1994), preliminary work suggests cosmological constraints from reconstructed voids at various redshifts are nonetheless possible. We plan to investigate these issues in future work.
Independent real-space distance information, combined with redshift data, yield the radial component of peculiar motion. In tandem, these measurement provide meaningful cosmological constraints (Qin et al., 2023; Hollinger & Hudson, 2024). Recent peculiar velocities surveys (Tully et al., 2016; Hong et al., 2019; Howlett et al., 2022) provide estimates for the nearby, low-redshift universe. Upcoming surveys (Saulder et al., 2023; Taylor et al., 2023; Mould et al., 2024; Said et al., 2024) offer a deeper view of the cosmic velocity field, including the interior of large voids.
@@ 7 Conclusion
Cosmic voids hold a bounty of information about the overall content, structure, and evolution of the universe (Alcock & Paczynski, 1979; Kirshner et al., 1981; Bertschinger, 1985; White et al., 1987; Regos & Geller, 1991; Blumenthal et al., 1992; Dubinski et al., 1993; Dekel & Rees, 1994; Lee & Park, 2009; Biswas et al., 2010). Because voids contain comparatively little matter, the effects of dark energy on the dynamics of visible objects lead, in principle, to testable model-dependent predictions (Colberg et al., 2005; Lee & Park, 2009; Tikhonov & Klypin, 2009; Maeda et al., 2011; Bos et al., 2012). The mere presence of the “largest void,” is one test of models (Blumenthal et al., 1992).
We explore how void dynamics can constrain the parameters of the underlying cosmology. We enhance the approaches of (Blumenthal et al., 1992) and (Schuster et al., 2023) where simple models and linear theory are applied to voids. We also build numerical infrastructure to simulate idealized voids and to generate $n$-body realizations of voids that emerge in a cosmological context. We focus on voids with sizes larger than 20 $h^{-1}$pc. We use simulation volumes that are between 250 $h^{-1}$pc and 5000 $h^{-1}$pc to a side. At the largest scale, an individual $n$-body realization provides representative samples of voids with radii up to $\sim$250 $h^{-1}$pc.
The numerical data here include fifteen independent $512^{3}$-particle $n$-body realizations (figure 15) and one $n=1024^{3}$ lightcone snapshot (figure 11). Other data sets in this project include more than 100 realizations with $n=512^{3}$, a half-dozen $n=1024^{3}$ realizations and one run with 2048 bodies, spanning a range of redshifts, cosmological models (table 2) and random number seeds (table 3). The adhesion algorithm (Appendix A) is a distinctive feature of this approach because it is at least an order of magnitude faster than full $n$-body simulations (Weinberg & Gunn, 1990). The adhesion approximation offers the possibility of maximum-likelihood fitting to void data with realizations evaluated at unique sets of cosmological parameters instead of the analytical approximation in equation (30). Thus, we can include deep voids ($\varepsilon>0.5$) in parameter estimation even when though the analytical approximation breaks down.
@@
  • Current maps of large-scale structure are primarily redshift surveys. We discuss interpretation of the properties of the largest voids in redshift space. Voids appear deeper in redshift space than they are in real physical space (the Kaiser effect; (Kaiser, 1987)).
  • To forecast future parameter estimation from voids in redshift surveys, we briefly mention adopting real-space reconstruction algorithms that extract physical void depths and outflow rates from redshift survey data (Dekel & Rees, 1994; Nadathur et al., 2019b, 2020c, 2020b; Aubert et al., 2022; Woodfinden et al., 2022, 2023; Radinović et al., 2023). This avenue may eventually be promising at $z\sim 1$ where the number of voids that can be accessed observationally is large.
@@ Appendix A The adhesion approximation
The adhesion approximation is an algorithm designed for $n$-body representations of large-scale structure (Weinberg & Gunn, 1990). The algorithm’s designers provide the details and rationale. We outline the steps to create a single realization with the implementation we use.
@@
    The adhesion algorithm requires an integration over time, performed using the growth factor $D$ (equation (12)) at $z_{i}$ and $z_{f}$. We set up a for-loop of ten-ish identical steps of size $\Delta D$ running from $D_{i}=D(z_{i})$ to $D_{f}=D(z_{f})$. To create a survey with redshifts that vary with position (defined on a grid, below), make $D(z_{f})$ depend on the real-space grid location. This step is a new addition to the original algorithm.
@@
    Define a cubic, triply-periodic computational domain with length $L$ on a side and $n^{3}$ gridpoints. Because we call Fast Fourier Transforms (FFTs), optimal choices for $n$ are powers of two. Because the adhesion approximation is a translinear method, the recommended grid spacing is about 1 Mpc (Weinberg & Gunn, 1990).
@@
    Realize a primordial density field in the Fourier domain choosing its normalization at the starting redshift. Define a grid in Fourier space with spacing $dk=n/L$. Calculate wave vector amplitude $|k|$ and evaluate $P(k)$ for each gridpoint. (We select either the CAMB algorithm (Lewis et al., 2000b) or the Eisenstein-Hu formulae as in the nbodykit package). The relative density at a gridpoint in $k$-space is
    $\delta(k)=sqrt(P(k;z)/L^{3})\times(u+iv)\times n^{3}$ (42)
@@
    Set up three coordinate grids so that $x,y,z$ are grid positions in the forward (spatial) domain (e.g., x values run from $-L+dL/2$ to $L+dL/2$; np.meshgrid is helful).
@@
The results are the components $(x,y,z,v_{x},v_{y},v_{z})$ of displaced gridpoints constituting the phase-space coordinates of individual $n$-body particles. We can reconstruct the density field by interpolating the “deformed” positions back onto a regular grid with a linear (or higher-order) interpolator. The novel part of our implementation is a “lookback mode” that calculates spatially dependent redshifts within the computational volume thus mimicking a redshift survey lightcone (Kitaura et al., 2016). In this mode we calculate the comoving distance to individual points from the center of the computational volume (Hogg, 1999) and then derive the expected redshift as seen by an observer there.

### v4

@@ 6.1 Reconstruction methods and peculiar velocity surveys
The large-scale structure revealed in redshift surveys, though affected by line-of-sight peculiar motion of individual galaxies, provides information about the underlying mass distribution in real space. If this real-space distribution were known, then linear theory would predict the peculiar velocities and hence the redshift-space location of galaxies within it. This mapping, at least on scales above $O(10)$ $h^{-1}$pc where the theory applies, is invertible. The real-space mass distribution and peculiar velocity flows can be reconstructed from redshift data [135, 136, 137, 138, 139, 140, 141, 142, 143, 144, 145, 146, 147, 148, 92, 149, 150, 151, 152, 153, 154, 155]. Recent work has extended reconstruction methods to include smaller scales where nonlinear dynamics affect redshift-space distortions [74, 156, 157, 158, 159, 160, 161].
Density reconstruction methods have been designed for the study of voids [130, 93]. Because the details of void selection are critical in measures of the void-galaxy correlator, reconstruction of the real-space density field prior to void identification is important so that redshift-space distortions do not introduce adverse selection effects [45, 96].
Reconstruction methods typically depend on assumptions about the background cosmology [106, 136] — the Hubble expansion rate $H$ and velocity growth factor $f$ are often required parameters in a reconstruction application. Therefore, reconstructed density and velocity fields on their own are not well suited for parameter estimation. They return the cosmological information from which they were derived. Following the lead of [136], preliminary work suggests cosmological constraints from reconstructed voids at various redshifts are nonetheless possible. We plan to investigate these issues in future work.
Independent real-space distance information, combined with redshift data, yield the radial component of peculiar motion. In tandem, these measurement provide meaningful cosmological constraints [157, 162]. Recent peculiar velocities surveys [163, 164, 165] provide estimates for the nearby, low-redshift universe. Upcoming surveys [166, 167, 168, 169] offer a deeper view of the cosmic velocity field, including the interior of large voids.
@@ 7 Conclusion
Cosmic voids hold a bounty of information about the overall content, structure, and evolution of the universe [14, 170, 10, 11, 12, 19, 171, 136, 95, 172]. Because voids contain comparatively little matter, the effects of dark energy on the dynamics of visible objects lead, in principle, to testable model-dependent predictions [173, 95, 174, 175, 176]. The mere presence of the “largest void,” is one test of models [19].
We explore how void dynamics can constrain the parameters of the underlying cosmology. We enhance the approaches of [19] and [70] where simple models and linear theory are applied to voids. We also build numerical infrastructure to simulate idealized voids and to generate $n$-body realizations of voids that emerge in a cosmological context. We focus on voids with sizes larger than 20 $h^{-1}$pc. We use simulation volumes that are between 250 $h^{-1}$pc and 5000 $h^{-1}$pc to a side. At the largest scale, an individual $n$-body realization provides representative samples of voids with radii up to $\sim$250 $h^{-1}$pc.
The numerical data here include fifteen independent $512^{3}$-particle $n$-body realizations (figure 15) and one $n=1024^{3}$ lightcone snapshot (figure 11). Other data sets in this project include more than 100 realizations with $n=512^{3}$, a half-dozen $n=1024^{3}$ realizations and one run with 2048 bodies, spanning a range of redshifts, cosmological models (table 2) and random number seeds (table 3). The adhesion algorithm (Appendix A) is a distinctive feature of this approach because it is at least an order of magnitude faster than full $n$-body simulations [94]. The adhesion approximation offers the possibility of maximum-likelihood fitting to void data with realizations evaluated at unique sets of cosmological parameters instead of the analytical approximation in equation (3.16). Thus, we can include deep voids ($\varepsilon>0.5$) in parameter estimation even when though the analytical approximation breaks down.
@@
  • Our numerical tracking of idealized void evolution identifies a simple approximation for the nonlinear relationship between the outflow rate ${H_{\text{eff}}}/H$ and the void depth $\varepsilon$ (equation (3.16)). This computationally efficient and surprisingly accurate formula enables precise estimation of cosmological model parameters from void properties in shell simulations and in large $n$-body realizations.
@@
  • Current maps of large-scale structure are primarily redshift surveys. We discuss interpretation of the properties of the largest voids in redshift space. Voids appear deeper in redshift space than they are in real physical space (the Kaiser effect; [106]).
  • To forecast future parameter estimation from voids in redshift surveys, we briefly mention adopting real-space reconstruction algorithms that extract physical void depths and outflow rates from redshift survey data [136, 78, 79, 132, 80, 97, 134, 49]. This avenue may eventually be promising at $z\sim 1$ where the number of voids that can be accessed observationally is large.
@@ Appendix A The adhesion approximation
The adhesion approximation is an algorithm designed for $n$-body representations of large-scale structure [94]. The algorithm’s designers provide the details and rationale. We outline the steps to create a single realization with the implementation we use.
@@
    The adhesion algorithm requires an integration over time, performed using the growth factor $D$ (equation (2.12)) at $z_{i}$ and $z_{f}$. We set up a for-loop of ten-ish identical steps of size $\Delta D$ running from $D_{i}=D(z_{i})$ to $D_{f}=D(z_{f})$. To create a survey with redshifts that vary with position (defined on a grid, below), make $D(z_{f})$ depend on the real-space grid location. This step is a new addition to the original algorithm.
@@
    Define a cubic, triply-periodic computational domain with length $L$ on a side and $n^{3}$ gridpoints. Because we call Fast Fourier Transforms (FFTs), optimal choices for $n$ are powers of two. Because the adhesion approximation is a translinear method, the recommended ( [94]) grid spacing is about 1 Mpc.
@@
    Realize a primordial density field in the Fourier domain choosing its normalization at the starting redshift. Define a grid in Fourier space with spacing $dk=n/L$. Calculate wave vector amplitude $|k|$ and evaluate $P(k)$ for each gridpoint. (We select either the CAMB algorithm [177] or the Eisenstein-Hu formulae as in the nbodykit package). The relative density at a gridpoint in $k$-space is
    $\delta(k)=\sqrt{(}P(k;z)/L^{3})\times(u+iv)\times n^{3}$ (A.1)
@@
    Set up three coordinate grids so that $x,y,z$ are grid positions in the forward (spatial) domain (e.g., x values run from $-L+dL/2$ to $L+dL/2$; np.meshgrid is helpful).
@@
The results are the components $(x,y,z,v_{x},v_{y},v_{z})$ of displaced gridpoints constituting the phase-space coordinates of individual $n$-body particles. We can reconstruct the density field by interpolating the “deformed” positions back onto a regular grid with a linear (or higher-order) interpolator. The novel part of our implementation is a “lookback mode” that calculates spatially dependent redshifts within the computational volume thus mimicking a redshift survey lightcone as in [178]. In this mode we calculate the comoving distance to individual points from the center of the computational volume [121] and then derive the expected redshift as seen by an observer there.